OpenStudy (anonymous):

In a binomial model, if p = .27, what is the value of q? A. .27 B. .63 C. .73 D. 1.0 Im getting C, but not sure if its the right answer.

OpenStudy (agent0smith):

q = 1-p = 1-0.27

OpenStudy (agent0smith):

You are correct as usual :)